OpenStudy (anonymous):

1. Halfway through the hockey season, a team has won 50 percent of the 40 games played. Their goal is to win 60 percent of their games. How many more games must they win? 2. On a 60- question test, Andre answered all the questions. He had 50 percent more questions correct than he had wrong. How many questions did he have correct?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

1. Let examine the data we have. Team has already played 40 games and it is half-season. That means the season has 40*2=80 games. Their goal is to win 60% of the games through the whole season. That means, they have to win 60% of 80 which is 0.6*80= 48 games. They already won 50% of 40 games played, meaning they won 0.5*40 = 20 games. If they won 20 games already, and need to win a total of 48 games, you can calculate how much more games they need to win;)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

how about 2

OpenStudy (anonymous):

2. Let's mark the correct answer with x and incorrect answers with y. Andre marked all questions (there are no unanswered, only correct and incorrect), so our equation is: x+y = 60 One more piece of information we know is the ratio between correct and incorrect answers. Andre had 50% more correct then incorrect: y = x + 0.5x y = 1.5x Now we combine the two equations: x + y = 60 x = 1.5y We insert the second equation in the first and we get: 1.5y + y = 60 2.5y = 60 From here we will get y (incorrect questions). Again, that value insert in one of the equations and we get x (correct questions). Hope this helps;) Feel free to ask if you don't understand something.

OpenStudy (anonymous):

thanks so much! you are like my new tutor! I go to tutoring, but it doesn't really help, and I have so much stuff that I need help with! im mostly bad with problem solving. I did everything in my book that I understood, but there's lots of things that I don't understand! thanks for being so kind and nice to me! this must be really easy for you, right? im in grade 8( going to grade8). thanks a lot!